In a hydraulic press, the area of the small piston is 25.0 cm2 and the area of the large piston is 325 cm2. If a force of 75.0 N is exerted on the small piston, find the weight of the object that can be lifted by the large piston.
1
Expert's answer
2020-03-30T08:07:26-0400

We use Pascal’s principle "\frac {F_1}{A_1}=\frac {F_2}{A_2}; 
F_2=F_1\times\frac{A_2}{A_1};
F_2=75 N \times \frac{325 {cm^2}}{25 {cm^2}};
F_2=975 N;

Answer: 975 N"